Week 7 - Midweek Update

Published on: Thu Jul 21 2011

This past week at the LTC has been quite productive. On Wednesday, Marty helped Gaby set up the AOM, and I was able to experiment with it. Before setting up the AOM I hooked up the oscillator to the oscilloscope, and verified that it was actually capable of producing 80MHZ waves. It was able to do this, and that is what we are using to drive the AOM. The oscillator has quite precise controls, and we can adjust it to things like 81Mhz etc… without trouble. I want to make measurements of the 1st order angle vs frequency, and also, I want to make the first order frequency, then rotate the AOM to measure intensity of the first order vs the angle of rotation of the AOM.

The second very exciting event was setting up laser Sam’s scanning Fabry Perot Laser Sam helped set it up quite a lot. The general arrangement is shown in this schematic:

For the rest of the week and the beginning of next, I’ll be using the two set ups to make measurements of the laser. Learning about the FP, and how to scan for resonant frequencies is one of my goals, but to do that I need to figure how how to prevent the beam from displacing with shifted frequency. Professor Metcalf had some suggestions about this, when he suggested I look up hetrodyne.